OpenStudy (anonymous):

Can you help me solve this inequality?? –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (nirmalnema):

first solve the brackets on both the sides..

OpenStudy (anonymous):

I know the left one will be -2x + 6, but I have no idea how to solve 5 - (x +3)

OpenStudy (anonymous):

it is 5-x-3

OpenStudy (anonymous):

no it would be 2-x

OpenStudy (nirmalnema):

2-x

OpenStudy (anonymous):

Can you explain @giligails

OpenStudy (anonymous):

x<=4

OpenStudy (nirmalnema):

-2x+6>= 2-x transfer the like terms at one side... u will get.. 6-2>=2x-x so x<=4

OpenStudy (anonymous):

@Falco276

OpenStudy (nirmalnema):

@StudyGirl21 did u understand

OpenStudy (anonymous):

5-(x+3)=5-x-3 5-3=2 thus the right side expression is 2-x

OpenStudy (falco276):

Oh gods, Im not that good with Inequalities. Sorry. :(

OpenStudy (anonymous):

No I'm still a little lost.

OpenStudy (falco276):

:(

OpenStudy (anonymous):

That's OK Falco.

OpenStudy (falco276):

yw. :)

OpenStudy (nirmalnema):

in which step u have confusion

OpenStudy (anonymous):

Can you show me how to work out the entire problem from the start @nirmalnema

OpenStudy (nirmalnema):

ok......

OpenStudy (nirmalnema):

first simplify the expression on both the side of inequality after that u will get///..... -2x+6>=2-x now u have to transfer the like terms at one side to further simplify it...

OpenStudy (anonymous):

from left and right expressions we get 6-2x>=2-x adding x on both sides 6-2x+x>=2-x+x 6-x>=2 subtracting 2 from both sides 6-x-2>=2-2 4-x>=0 adding x on both sides 4-x+x>=x x<=4

OpenStudy (nirmalnema):

do u know how to do that..

OpenStudy (anonymous):

thanks giligails, I'm a little slow. :)

OpenStudy (nirmalnema):

after that u will get.. 6-2>=2x-x so x<=4